The Golus Elimination method: Example: Solve the system: 2x +y +2 = 5 4x-by = -2 -2x + 7y +22=9
Answer #1

The Gauss elimination method Given system is 2x + y +2=5 4n-by= -28+7y +27 = 9 the system is of the form The Augmented matrix